# Distributions and Direct Parametrization for Stable Stochastic State-Space Models

### v1

@@ VI Conclusion and Future Work
In this work, we have introduced a direct parametrization for continuous-time stochastic state-space models that guarantees external stability via the stochastic bounded-real lemma. This framework not only enables the construction of probabilistic priors that enforce stability almost surely but also facilitates robust Bayesian inference, as demonstrated by our simulation study.
Future work will explore extensions in several directions. One avenue is to incorporate alternative stability notions and extend the approach to linear parameter-varying systems. Another direction is to develop more efficient Bayesian inference methods, for example by leveraging variational inference techniques.

### v2

@@ VI Conclusion and Future Work
In this work, we have introduced a direct parametrization for continuous-time stochastic SSMs that guarantees external stability via the stochastic BRL. This framework not only enables the construction of probabilistic priors that enforce stability almost surely but also facilitates robust Bayesian inference, as demonstrated by our simulation study.
Future work will explore incorporating alternative stability notions and extending the approach to linear parameter-varying systems. Another direction is to experiment with and develop more efficient Bayesian inference methods for high-dimensional SSMs, for example, by leveraging variational inference techniques.